OpenStudy (anonymous):

Simplify following sq. root problems. 5 sq. root 27 + 7 sq. root 3

OpenStudy (a_clan):

5 sq. root 27 + 7 sq. root 3 \[5\sqrt{3*3*3}+7\sqrt{3}\] \[5*3*\sqrt{3}+7\sqrt{3}\] \[15\sqrt{3}+7\sqrt{3}\] \[22\sqrt{3}\]